Grace is 8 years older than her brother Sam. The sum of their ages is 24. How old is Grace?

Answer by lefty4ever26(59)

Need to make 2 equations.

G=8+S
G+S=24 Substitute G=8+S into second equation.
G+S=24
(8+S)+S=24 Solve for S.
8+2S=24
2S=16
S=8 So we know that Sam is 8 yrs old. Substitute S back into the first equation.
G=8+S
G=8+8
G=16 So Grace is 16 yrs old.